WACC.  Eric has another​ get-rich-quick idea, but needs funding to support it. He chooses an​ all-debt funding scenario. He will borrow ​$6,463 from​ Wendy, who will charge him 5​% on the loan. He will also borrow ​$5,446 from​ Bebe, who will charge him 7​% on the​ loan, and ​$2,091 from​ Shelly, who will charge him 13​% on the loan. What is the weighted average cost of capital for​ Eric?
Answer #1
Total Loan portfolio value = Value of loan from Wendy + Value of loan from Bebe + Value of loan from Shelly
=6463+5446+2091
=14000
Weight of loan from Wendy = Value of loan from Wendy/Total Loan portfolio Value
= 6463/14000
=0.4616
Weight of loan from Bebe = Value of loan from Bebe/Total Loan portfolio Value
= 5446/14000
=0.389
Weight of loan from Shelly = Value of loan from Shelly/Total Loan portfolio Value
= 2091/14000
=0.1494
Cost of capital of Loan portfolio = Weight of loan from Wendy*Cost of capital of loan from Wendy+Weight of loan from Bebe*Cost of capital of loan from Bebe+Weight of loan from Shelly*Cost of capital of loan from Shelly
Cost of capital of Loan portfolio = 5*0.4616+7*0.389+13*0.1494
Cost of capital of Loan portfolio = 6.9729